OpenStudy (kewlgeek555):

Help with Two Algebra Questions About Solving Systems of Equation Algebraically? Please?!

OpenStudy (kewlgeek555):

@ganeshie8 @DebbieG @mathstudent55 @agent0smith @nincompoop @ranga

OpenStudy (kewlgeek555):

Choose the equivalent system of linear equations that will produce the same solution as the one given below. 6x + 2y = -6 3x - 4y = -18 a. 12x + 4y = -12 15x = -30 b. 8x + 4y = -4 14x = -10 c. 6x - 8y = -36 -6y = -42 d. 6x - y = -15 3y = 9

OpenStudy (kewlgeek555):

I think that it is either a or c...

OpenStudy (anonymous):

have you tried solving the system of equations given to find x and y, then plugging in x and y into the choices?

OpenStudy (anonymous):

hint: they are using elimination to get the second equation in the answers. why don't you give it a shot with the original eqns...

OpenStudy (anonymous):

the person isn't even here!

OpenStudy (anonymous):

:/

OpenStudy (kewlgeek555):

I'm right here. It's just that OpenStudy is not working and is not giving notifications.

OpenStudy (kewlgeek555):

I can't find x and y @DemolisionWolf T.T

OpenStudy (anonymous):

did you eliminate?

OpenStudy (anonymous):

If your not sure how to find x and y using substitution, then you will want to use elimination to solve this problem like pgpilot326 says

OpenStudy (kewlgeek555):

Well, I got the answer - it was a.

OpenStudy (kewlgeek555):

thnaks.

OpenStudy (anonymous):

ur welcome sorry it took so long to get back here, notifications are't working